# Secure and Efficient Access Control Framework for Computer-Use Agents via Context Space

### v3

@@ 4.2. Intent-aware Context Space
where ctx_idctx\_id is the unique identifier of the context. constraintconstraint is a logical expression that must evaluate to true against the corresponding context value for safe execution. Constraints are typically expressed in the form val1 operator val2, where val1 refers to the current context value, and val2 is either a constant or the value of another context. operator is a relational symbol like =, <, $\in$, and $\subset$. guidanceguidance provides human-readable hints when constraint validation fails, helping users understand why the function was blocked. To centrally manage all contexts within a context space and facilitate policy validation at runtime, we introduce the concept of context vector. The vector serves as a unified data structure that maintains the current runtime values of all relevant contexts defined in the policies. At runtime, the current state is represented by a context vector $CV$:
@@
where $validate(rule.constraint,CV)$ returns true if the constraint condition is satisfied by the corresponding context value in the current context vector $CV$.
@@ 4.3. LLM-based Context Analyzer
Challenges in GUI analysis. For API- and CLI-based apps, policy generation is straightforward due to the availability of well-defined function specifications. Specifically, API tools usually provide detailed documentation on purpose, parameters, and behavior (qin2023toolllm; website:langchaintool), while CLIs usually offer manuals specifying command functionality and usage (website:linuxman; website:powershelldoc), which enables LLMs to generate high-quality context spaces with sufficient semantic information on function behavior. However, GUI agents present greater challenges. Concretely, GUI agents interact with apps through generic interaction primitives (such as screen capture, clicking, scrolling, and text input) rather than explicit functional interfaces. Even worse, GUI apps typically lack functional documentation. While human users understand app usage through interaction, this knowledge is rarely formalized in machine-readable specifications, making it difficult to identify the security-sensitive functions that require protection. As a result, no existing agent protection approach has demonstrated automated end-to-end policy generation for GUI apps. Existing approaches for GUI functionality discovery include static program analysis and GUI exploration tools. However, our experiments (§7.2) reveal limitations in both approaches: static analysis often fails to capture the complete scope of GUI functionalities due to dynamic binding and complex UI frameworks, while GUI exploration tools suffer from scalability issues and frequently encounter infinite loops or state explosion problems that prevent comprehensive coverage.
@@ 4.4. Optimized Context Manager
Challenges in performance. The context manager faces two primary performance challenges. First, LLM-based context extraction from user instructions introduces additional inference latency. One simple approach is to request the LLM to analyze the relevant intent and extract context values after the agent determines the function to perform, which would introduce inference latency for every agent action, severely degrading performance (shi2025progent; lee2025verisafe). Second, context management for complex apps creates substantial computational overhead. Large-scale apps, particularly GUI apps, often have extensive context spaces involving numerous contexts. Frequent acquisition and updates of all context values incur unnecessary computational costs, increased resource consumption, and additional energy drain. Moreover, loading these context spaces is time-consuming, especially when switching between multiple apps, resulting in a poor user experience.
Our solution. Figure 5 shows how we address these challenges by adopting thorough optimizations through parallel processing and systematic context management. To mitigate LLM inference overhead, we employ a concurrent approach for extracting user intents and related contexts from user instructions. When users send requests, the instruction is simultaneously dispatched to both the agent and a dedicated LLM-based intent extractor in the CSAgent service. The manager gathers all function intents and parameter contexts from the context space and passes them to the extractor. The extractor then analyzes the user instruction and selects all relevant intents and extracts contexts that correspond to the current request. Additionally, for function intents, we prompt the LLM to infer reasonable execution sequences that can be used to detect order violations and prevent errors caused by out-of-sequence actions. When the agent determines the next function, we use the selected intent to retrieve the policy. However, LLMs’ selections are not always accurate. To refine the decision, we use cosine similarity between user instructions and intent descriptions: for missing selections, we apply a threshold to determine if the instruction truly relates to the function; for multiple selections, we choose the intent with the highest similarity score.
@@ 4.5. Policy Evolution Framework
Challenges in policy reliability. We acknowledge that LLMs’ uncertainty makes it difficult to guarantee policy correctness and completeness. CSAgent relies on LLMs to generate context space (including intents, contexts and policies), analyze GUI event handlers, and select intent during policy retrieval. Inaccuracies and omissions in their content can compromise access control effectiveness. Since writing completely correct and comprehensive policies is impractical (akin to bug-free software), the only viable mitigation is continuous policy refinement. Manual policy review is labor-intensive, requiring domain experts with deep knowledge of both app functionality and security.
Our solution. Static policy offers a key advantage of consistency across deployments, i.e., all users use identical policies, and we can centralize policy maintenance. Building on this, we propose the Policy Evolution Framework (PEF), a tool that facilitates policy refinement through app update synchronization and runtime feedback analysis, as shown in Figure 4 (orange arrows). When an app undergoes feature changes, developers can provide update information (e.g., patches or changelogs) to PEF. The framework automatically updates the context space. For new functions, PEF uses the context analyzer to generate fresh policies. For modified or removed functions, PEF locates the corresponding policies and either regenerates or removes them accordingly. For runtime feedback analysis, the CSAgent service continuously collects policy-related events through logging. Policy anomalies are identified through two mechanisms. During policy retrieval, the service logs cases where target functions or intents cannot be located, capturing user instructions and missing elements. During policy validation, the service records complete validation sequences, including policies and their outcomes, context values, and user instructions. When policy validation fails, the system requests user confirmation before blocking. If users indicate the function should proceed, this suggests potential policy deficiencies. Conversely, when validation succeeds but users believe the function should be blocked, they can actively report anomalies. PEF collects these logs, analyzes them through statistics and LLMs, and provides policy improvement suggestions. Developers can directly adopt these suggestions or manually intervene as needed. This approach mirrors existing OS crash reporting mechanisms, with similar user privacy controls allowing opt-out from logging and data collection.
@@ 5. Implementation
Automatic policy generation and evolution. As outlined in §4.3, we use LLMs (DeepSeek-R1 (guo2025deepseek)) for policy generation and refinement. For API and CLI agents, the LLM analyzes function documentation to generate policies. For GUI agents, we target Android apps since Android is widely used and open-source. For a given app’s source code, our context analyzer first uses text matching to identify files related to GUI operations, then provides these files to the LLM for analysis of event handlers. The LLM outputs results in JSON format, including GUI element details like class and resource ID, and methods invoked by the corresponding handlers. We then use static analysis tools (CodeQL (website:codeql)) to construct call graphs. For policy evolution, we implement a runtime logging system that captures task information, validation events, user feedback, and execution traces. During refinement, the LLM analyzes the logs alongside the existing context space and update history to identify potential utility and security issues, then systematically updates policies to address these deficiencies.
Policy retrieval. Before each agent action, the CSAgent service retrieves the corresponding policy via function and intent indexing. For API/CLI agents, the mapping is direct, with each tool call or CLI command corresponding to a function entry. GUI agents present a more complex scenario, as their actions consist of GUI controls rather than function calls. To address this, CSAgent captures the screen’s GUI tree and uses the agent’s target coordinates to identify the intended GUI element, extracting its package, class, and resource ID attributes to map to the corresponding function.
Context acquisition and policy validation. Context values come from user instructions, device systems, and agent operations. During intent selection, the intent extractor LLM parses user requests and environmental data (e.g., configurations) to obtain contexts. Agent-related contexts are captured in two ways: parameter contexts are obtained when intercepting agent actions, while history contexts are updated after action completion. For system-level contexts, we currently support CLI-based acquisition via tools like ADB (website:adb). Another approach is implementing dedicated system services to provide unified interfaces for context acquisition, which requires platform vendor participation in the ecosystem.
@@ 6.4. Threat Analysis
Policy reliability. The security improvement provided by CSAgent largely depends on the correctness and completeness of the policies. While generating fully correct and complete policies is unrealistic, CSAgent mitigates this issue by leveraging more powerful reasoning models to generate higher-quality policies. Additionally, we have fallback mechanisms to address situations where policies are incomplete (e.g., missing intents or functions). In such cases, the system requests user intervention or prompts the agent to re-plan, ensuring that uncertain actions are blocked.
@@ 7.1. Experimental Setup and Benchmarks (RQ1)
To evaluate CSAgent across different agent interaction modalities, we employ three benchmarks: AgentBench  (liu2023agentbenchevaluatingllmsagents) (OS environment), AgentDojo  (agentdojo), and AndroidWorld  (rawles2024androidworld), which respectively evaluate CLI-based, API-based, and GUI-based agents. We employ DeepSeek-V3  (liu2024deepseekv3) as our runtime text model, and Seed1.6  (website:seed16) as the multimodal model. The integration process with these benchmarks closely mirrors real-world CSAgent deployment, consisting of two steps. First, we analyze the applications within each benchmark, including shell commands, tools, and apps, to generate context spaces using our context analyzer. Second, we adapt the agent frameworks to the CSAgent service with minimal code modifications: establishing an RPC connection, invoking the intent extraction method upon user request arrival, and calling validation and context update methods before and after agent actions. Our successful support of these benchmarks demonstrates CSAgent’s strong compatibility (satisfying G4).
Our evaluation methodology encompasses three primary dimensions across all benchmarks: utility, security, and overhead. Utility measures the agent’s task completion capability under CSAgent protection, security evaluates CSAgent’s ability to block unintended behaviors, and overhead quantifies the latency and token consumption introduced by CSAgent. For AndroidWorld specifically, we also conduct an analysis of CSAgent’s code identification capabilities for context space generation. We compare it against UI-CTX  (li2025uictx) and AutoDroid  (wen2024autodroid), representing state-of-the-art approaches in static analysis and GUI exploration, respectively. For security evaluation, all the context policies are generated by the context analyzer (satisfying G3). While AgentDojo provides built-in adversarial tests, we construct additional security test cases for AgentBench and AndroidWorld. These test cases consist of mismatched user requests and agent actions (i.e., simulating agent actions that violate the user intent), creating a series of pairs that serve as input to evaluate capability to detect anomalous agent behaviors that deviate from user intentions. Notably, the policies we test are generated solely based on the apps themselves, and the policy generation process has no access to the test set.
To emphasize effectiveness, we design our experiments around CSAgent and two other agent configurations. First, we evaluate a Vanilla Agent that operates without any security protection mechanisms, serving as our analysis baseline. Second, we implement a Pre-execution Validation Agent (PVAgent) that employs dynamic policy generation and validation before executing each action (similar to (tsai2025contextualagent; shi2025progent)). Finally, we evaluate CSAgent-RF, a refined variant of CSAgent that leverages PEF, to further demonstrate its effectiveness.
@@ 7.2. Capability of GUI Analysis (RQ2)
Figure 6 demonstrates CSAgent’s capability in extracting GUI elements and event handlers from open-source apps and AOSP system apps used by AndroidWorld. UI-CTX was configured with default settings, while AutoDroid employed the deep learning-based strategy Humanoid (li2020humanoiddeeplearningbasedapproach). The results show that CSAgent extracts 0.93× more GUI element information (including handlers) than AutoDroid and 3.12× more than UI-CTX (satisfying G3). While some GUI elements are identified by these two baseline methods but missed by CSAgent, most do not produce actual control effects (e.g., merely opening menus rather than clicking specific options). We prompt our context analyzer to ignore these non-functional elements to optimize token consumption. We did not compare to source code-based analysis primarily due to the lack of available tools. Our approach differs by using LLMs to identify GUI-related code, offering better generalization capabilities. Importantly, CSAgent can complement other methods to achieve more comprehensive coverage.
@@ 7.3. Effectiveness of CSAgent Access Control (RQ3)
The reduction in UUA is primarily because many tasks in AgentDojo depend on contexts from untrusted information sources. In our experiments, we configured local information sources (such as file systems and calendars) as trusted, while online sources (e.g., web and email) as untrusted. Consequently, CSAgent blocks execution of related actions and requests user confirmation, which is reasonable behavior. With more fine-grained trust distinctions for information sources, CSAgent’s utility could be further improved.
@@ 7.4. Performance and Cost Overhead (RQ4)
Figure 7 illustrates the utility drop and extra latency introduced by CSAgent. On all three benchmarks, both utility drop and latency overhead introduced by CSAgent are less than 5%, compared to PVAgent’s overhead of 58.57%, 34.28%, and 103.98% on AgentBench, AndroidWorld, and AgentDojo, respectively. Additionally, we found that 80.08% of contexts are classified as cold or warm temperature levels. These results demonstrate that our optimized context manager efficiently performs user intent extraction, context updates and policy verification (satisfying G2). For utility, CSAgent causes a 15.58% decrease on average. The degradation on AgentDojo is more pronounced, primarily due to task complexity and insufficient policy quality, as discussed earlier. Notably, upon a policy validation failure during evaluation, the agent is prompted to either explore alternative execution paths or terminate the task. In real deployments, user intervention or agent reflection could resume such tasks, further reducing the impact. To explore this potential, we evaluate a variant incorporating agent reflection upon validation failures. On AgentDojo, this approach reduces the utility drop to 9.93% and improves UUA to 51.79%. However, it incurs a 21.04% latency overhead and increases ASR to 0.455%, as additional LLM requests introduce both computational costs and increased uncertainty. We also evaluated CSAgent’s impact on task completion steps. Overall, the impact is minimal, indicating that CSAgent’s decisions to allow or block agent actions are relatively accurate.
((a))
((b))
@@
We further evaluated the complexity of context spaces and their loading latency. Across the three benchmarks (AgentBench, AgentDojo, and AndroidWorld), the average number of policies per application, which is also the number of predicted user intents, is 120, 13.25, and 150.4, respectively, with context counts of 187.0, 25.75, and 238.65. We also analyzed real-world complex apps from AndroidWorld to explore the relationship between app size and context space size. As shown in Figure LABEL:fig:n_policy, the policy count grows at a diminishing rate as app size increases. This is mainly because we generate policies only for security-critical functions, which are typically limited in number. Figure LABEL:fig:lat_perc shows the loading latency distribution, revealing that 96.00% of context spaces can be loaded within 1 seconds. Loading latency correlates directly with context space size and I/O latency. When context spaces are cached, retrieval latency becomes negligible. Therefore, beyond our LRU replacement policy, larger context spaces (such as that for the settings app) can be permanently fixed in cache to further optimize performance.
@@
Table 5 presents CSAgent’s per task token consumption compared to the vanilla agent. The geometric mean of consumption overhead is 35.30%, with over 96.41% consisting of prompt tokens, which are less expensive than completion tokens. Generally, more complex tasks have proportionally lower overhead. For instance, AgentBench tasks have low average token consumption, so the large context space results in higher overhead. Conversely, AndroidWorld tasks are highly complex, requiring multimodal capabilities for screen recognition, leading to very high task token consumption and lower overhead. In comparison, runtime policy generation methods introduce 1.61× additional token consumption on AgentDojo (shi2025progent). CSAgent’s low token consumption results from development-phase policy generation that applies across all devices and users without regeneration, and from requiring only one additional LLM inference per task regardless of complexity (satisfying G2).

### v4

@@ 4.2. Intent-aware Context Space
where ctx_idctx\_id is the unique identifier of the context. constraintconstraint is a logical expression that must evaluate to true against the corresponding context value for safe execution. Constraints are constructed using atomic predicates combined with logical connectives (e.g., $\land$, $\lor$, $\neg$). An atomic predicate typically takes the form val1 operator val2, where val1 refers to the current context value, and val2 represents a reference target. Crucially, val2 supports flexible assignment: it can be a pre-defined constant, a user-specified configuration, or a dynamic value derived from another context, ensuring that security policies adapt to individual usage patterns. The operator includes relational symbols (e.g., $=,<,\neq$) and set operations (e.g., $\in,\subset$). This flexible definition allows policies to express complex logic, such as ensuring a user role belongs to a permitted set ($role\in\{admin,owner\}$) or validating value ranges ($10<amount<500$). guidanceguidance provides human-readable hints when constraint validation fails, helping users understand why the function was blocked. To centrally manage all contexts within a context space and facilitate policy validation at runtime, we introduce the concept of context vector. The vector serves as a unified data structure that maintains the current runtime values of all relevant contexts defined in the policies. At runtime, the current state is represented by a context vector $CV$:
@@
where $validate(rule.constraint,CV)$ returns true if the constraint condition is satisfied by the corresponding context value in the current context vector $CV$. This validation logic enforces a default-deny security model: any sensitive function execution is blocked unless explicitly authorized by a matching policy and satisfied context constraints. While the policy enforces the conjunction (AND) of rules, disjunctive scenarios (OR) are supported through flexible constraints or distinct user intents for mutually exclusive contexts. This design flattens complex branching into linear paths, ensuring security (via default-deny) without sacrificing expressiveness.
@@ 4.3. LLM-based Context Analyzer
Challenges in GUI analysis. For API- and CLI-based apps, policy generation is straightforward due to the availability of well-defined function specifications. Specifically, API tools usually provide detailed documentation on purpose, parameters, and behavior (Qin et al., 2023; LangChain, 2025c), while CLIs usually offer manuals specifying command functionality and usage (Kerrisk, 2025; Microsoft, 2025b), which enables LLMs to generate high-quality context spaces with sufficient semantic information on function behavior. However, GUI agents present greater challenges. Concretely, GUI agents interact with apps through generic interaction primitives (such as screen capture, clicking, scrolling, and text input) rather than explicit functional interfaces. Even worse, GUI apps typically lack functional documentation. While human users understand app usage through interaction, this knowledge is rarely formalized in machine-readable specifications, making it difficult to identify the security-sensitive functions that require protection. As a result, no existing agent protection approach has demonstrated automated end-to-end policy generation for GUI apps. Existing approaches for GUI functionality discovery include static program analysis and GUI exploration tools. However, our experiments (§7.2) reveal limitations in both approaches: static analysis often fails to capture the complete scope of GUI functionalities due to dynamic binding and complex UI frameworks, while GUI exploration tools suffer from scalability issues and frequently encounter infinite loops or state explosion problems that prevent comprehensive coverage.
@@ 4.4. Optimized Context Manager
Challenges in performance. The context manager faces two primary performance challenges. First, LLM-based context extraction from user instructions introduces additional inference latency. One simple approach is to request the LLM to analyze the relevant intent and extract context values after the agent determines the function to perform, which would introduce inference latency for every agent action, severely degrading performance (Shi et al., 2025a; Lee et al., 2025). Second, context management for complex apps creates substantial computational overhead. Large-scale apps, particularly GUI apps, often have extensive context spaces involving numerous contexts. Frequent acquisition and updates of all context values incur unnecessary computational costs, increased resource consumption, and additional energy drain. Moreover, loading these context spaces is time-consuming, especially when switching between multiple apps, resulting in a poor user experience.
Our solution. Figure 5 shows how we address these challenges by adopting thorough optimizations through parallel processing and systematic context management. To mitigate LLM inference overhead, we employ a concurrent approach for extracting user intents and related contexts from user instructions. When users send requests, the instruction is simultaneously dispatched to both the agent and a dedicated LLM-based intent extractor in the CSAgent service. To improve accuracy and efficiency, we employ a coarse-to-fine retrieval strategy before invoking the LLM. Specifically, the manager first uses cosine similarity to efficiently filter relevant intent candidates from the context space based on the user instruction. These filtered candidates, along with parameter contexts, are then passed to the extractor. The extractor analyzes the user instruction to identify the valid intent and extract contexts that correspond to the current request. For function intents, we also prompt the LLM to infer reasonable execution sequences to prevent out-of-sequence errors. To address potential ambiguities where the LLM identifies multiple intents for a single function, we select the intent with the highest similarity score to ensure precise policy retrieval. When the agent determines the next function, we use this finalized intent to retrieve the policy.
@@ 4.5. Policy Evolution Framework
Challenges in policy reliability. We acknowledge that LLMs’ uncertainty makes it difficult to guarantee policy correctness and completeness. CSAgent relies on LLMs to generate context space (intents, contexts and policies), analyze GUI event handlers, and select intent during policy retrieval. Inaccuracies and omissions in their content can compromise access control effectiveness. Since writing completely correct and comprehensive policies is impractical (akin to bug-free software), the only viable mitigation is continuous policy refinement. Manual policy review is labor-intensive, requiring domain experts with deep knowledge of both app features and security.
Our solution. We treat policy engineering as analogous to software engineering: initial LLM-generated policies are ”beta code” that must undergo rigorous testing and iterative refinement. Static policy offers a key advantage of consistency across deployments: all users operate under identical policies, enabling centralized quality assurance. Building on this, we propose the Policy Evolution Framework (PEF), a comprehensive toolchain that ensures policy reliability through generation-time validation, app update synchronization, and runtime feedback-based evolution, as shown in Figure 4 (orange arrows). When a new context space is generated, PEF performs initial validation: it verifies the format correctness, then checks function coverage for API/CLI apps or validates the existence and consistency of UI element classes, identifiers, and event handlers. When an app undergoes feature changes, developers provide update information (e.g., patches or changelogs) to PEF, which automatically updates the context space. For new functions, PEF uses the context analyzer to generate fresh policies. For modified or removed functions, PEF locates the corresponding policies and either regenerates or removes them accordingly.
For runtime feedback analysis, the CSAgent service logs the full execution trace: user request, extracted intent, context evaluation process, and validation result. Policy anomalies are identified through two mechanisms. During policy retrieval, the service logs cases where target functions or intents cannot be located, capturing user instructions and missing elements. During policy validation, when validation fails, the system requests user confirmation before blocking; if users indicate the function should proceed, this suggests potential policy deficiencies. Conversely, when validation succeeds but users believe the function should be blocked, they can actively report anomalies. The CSAgent service filters anomalous logs during idle periods and uploads them to PEF for analysis. PEF then identifies the root cause by prompting the LLM with the runtime error log, the target function’s specification or code, and the failed policy, subsequently generating improvement suggestions for developers, who can directly adopt these suggestions or manually intervene as needed. This approach mirrors existing OS crash reporting mechanisms, with similar user privacy controls allowing opt-out from logging and data collection. To complement offline refinement, we support online parameter adaptation for user-customizable values. These parameters (e.g., threshold or recipient whitelists) are initialized with strict defaults (e.g., empty sets) to ensure safety. When a legitimate action is blocked due to these conservative constraints, the system prompts the user for verification. Upon approval, the specific context value is dynamically incorporated into the trusted set, relaxing the policy for future execution.
@@ 5. Implementation
Automatic policy generation and evolution. As outlined in §4.3, we use LLMs (DeepSeek-R1 (Guo et al., 2025)) for policy generation and refinement. For API and CLI agents, the LLM analyzes function documentation to generate policies. For GUI agents, we target Android apps since Android is widely used and open-source. For a given app’s source code, our context analyzer first uses text matching to identify files related to GUI operations, then provides these files to the LLM for analysis of event handlers. The LLM outputs results in JSON format, including GUI element details like class and resource ID, and methods invoked by the corresponding handlers. We then use static analysis tools (CodeQL (GitHub, 2021)) to construct call graphs. For policy evolution, we implement a runtime logging system that captures task information, validation events, user feedback, and execution traces. During refinement, the LLM analyzes the logs alongside the existing context space and update history to identify potential utility and security issues, then systematically updates policies to address these deficiencies.
Policy retrieval. Upon receiving a request, we first retrieve a set of relevant (Function,Intent) candidates from the context space using cosine similarity. Next, we feed the request, these candidates, and policy-defined contexts to the intent extractor LLM, which then identifies the valid intent and extracts specific context values, returning them in a list. Subsequently, before each agent action, the CSAgent service retrieves the corresponding policy via function and intent indexing. For API/CLI agents, the mapping is direct, with each tool call or CLI command corresponding to a function entry. GUI agents present a more complex scenario, as their actions consist of GUI controls rather than function calls. To address this, CSAgent captures the screen’s GUI tree and uses the agent’s target coordinates to identify the intended GUI element, extracting its package, class, and resource ID attributes to map to the corresponding function.
Context acquisition and policy validation. Context values come from user instructions, systems, and agent operations. During intent selection, the intent extractor LLM parses user requests and environmental data (e.g., configurations) to obtain contexts. Agent-related contexts are captured in two ways: parameter contexts are obtained when intercepting agent actions, while history contexts are updated after action completion. For system state contexts, values are acquired through CLI (e.g., shell or ADB (Google, 2025b)) or APIs accessible to the agent. Another approach is implementing dedicated system services to provide unified interfaces for context acquisition, which requires platform vendor participation in the ecosystem.
@@ 6.4. Threat Analysis
Policy reliability. The security improvement provided by CSAgent largely depends on the correctness and completeness of the policies. While generating fully correct and complete policies is unrealistic, CSAgent mitigates this issue by leveraging more powerful reasoning models to generate higher-quality policies and iteratively improve them through the PEF. Additionally, we have fallback mechanisms to address situations where policies are incomplete (e.g., missing intents or functions). In such cases, the system requests user intervention or prompts the agent to re-plan, ensuring that uncertain actions are blocked.
@@ 7.1. Experimental Setup and Benchmarks (RQ1)
To evaluate CSAgent across diverse interaction modalities, we employ three benchmarks: AgentBench (Liu et al., 2023) (CLI), AgentDojo (Debenedetti et al., 2024) (API), and AndroidWorld (Rawles et al., 2024) (GUI). We use DeepSeek-V3 (Liu et al., 2024a) and Seed1.6 (ByteDance, 2025) as the runtime text and multimodal models, respectively. The integration mirrors real-world deployment in two steps: first, we generate context spaces for target tools/apps using our analyzer; second, we adapt agent frameworks to the CSAgent service via minimal RPC hooks for intent extraction, validation, and context updates. This successful integration demonstrates CSAgent’s strong compatibility (G4).
Our evaluation methodology encompasses three primary dimensions across all benchmarks: utility, security, and overhead. Utility measures the agent’s task completion capability under CSAgent protection, security evaluates CSAgent’s ability to block unintended behaviors, and overhead quantifies the latency and token consumption introduced by CSAgent. For AndroidWorld specifically, we also conduct an analysis of CSAgent’s code identification capabilities for context space generation. We compare it against UI-CTX  (Li et al., 2025b) and AutoDroid  (Wen et al., 2024), representing state-of-the-art approaches in static analysis and GUI exploration, respectively. For security evaluation, all the context policies are generated by the context analyzer (satisfying G3). While AgentDojo provides built-in adversarial tests, we construct additional security test cases for AgentBench and AndroidWorld. These test cases consist of mismatched user requests and agent actions (i.e., simulating agent actions that violate the user intent), creating a series of pairs that serve as input to evaluate capability to detect anomalous agent behaviors that deviate from user intentions. Notably, the policies we test are generated solely based on the apps themselves, and the policy generation process has no access to the test set.
To emphasize effectiveness, we design our experiments around CSAgent and two other agent configurations. First, we evaluate a Vanilla Agent that operates without any security protection mechanisms, serving as our analysis baseline. Second, we implement a Pre-execution Validation Agent (PVAgent) that employs dynamic policy generation and validation before executing each action (similar to (Tsai and Bagdasarian, 2025; Shi et al., 2025a)). Finally, we evaluate CSAgent-RF, a refined variant of CSAgent that leverages PEF, to further demonstrate its effectiveness.
@@ 7.2. Capability of GUI Analysis (RQ2)
Figure 6 demonstrates CSAgent’s capability in extracting GUI elements and event handlers from open-source apps and AOSP system apps used by AndroidWorld. UI-CTX was configured with default settings, while AutoDroid employed the deep learning-based strategy Humanoid (Li et al., 2019). The results show that CSAgent extracts 0.93× more GUI element information (including handlers) than AutoDroid and 3.12× more than UI-CTX (satisfying G3). While some GUI elements are identified by these two baseline methods but missed by CSAgent, most do not produce actual control effects (e.g., merely opening menus rather than clicking specific options). We prompt our context analyzer to ignore these non-functional elements to optimize token consumption. We did not compare to source code-based analysis primarily due to the lack of available tools. Our approach differs by using LLMs to identify GUI-related code, offering better generalization capabilities. Importantly, CSAgent can complement other methods to achieve more comprehensive coverage.
@@ 7.4. Performance and Cost Overhead (RQ4)
Figure 7 illustrates the utility drop and extra latency introduced by CSAgent. On all three benchmarks, latency overhead introduced by CSAgent is less than 5%, compared to PVAgent’s overhead of 58.57%, 34.28%, and 103.98% on AgentBench, AndroidWorld, and AgentDojo, respectively. Additionally, we found that 80.08% of contexts are classified as cold or warm temperature levels. These results demonstrate that our optimized context manager efficiently performs user intent extraction, context updates and policy verification (satisfying G2). For utility, CSAgent causes a 15.58% decrease on average, significantly outperforming PVAgent. The degradation on AgentBench and AndroidWorld is less than 5%, but that on AgentDojo is more pronounced, primarily due to task complexity and insufficient policy quality, as discussed earlier. Notably, upon a policy validation failure during evaluation, the agent is prompted to either explore alternative execution paths or terminate the task. In real deployments, user intervention or agent reflection could resume such tasks, further reducing the impact. To explore this potential, we evaluate a variant incorporating agent reflection upon validation failures. On AgentDojo, this approach reduces the utility drop to 9.93% and improves UUA to 51.79%. However, it incurs a 21.04% latency overhead and increases ASR to 0.455%, as additional LLM requests introduce both computational costs and increased uncertainty. We also evaluated CSAgent’s impact on task completion steps. Overall, the impact is minimal, indicating that CSAgent’s decisions to allow or block agent actions are relatively accurate.
@@
We further evaluated the complexity of context spaces and their loading latency. Across the three benchmarks (AgentBench, AgentDojo, and AndroidWorld), the average number of policies per application, which is also the number of predicted user intents, is 120, 13.25, and 150.4, respectively, with average context counts of 187.0, 25.75, and 238.65. We also analyzed real-world complex apps from AndroidWorld to explore the relationship between app size and context space size. As shown in Figure LABEL:fig:n_policy, the policy count grows at a diminishing rate as app size increases. This is mainly because we generate policies only for security-critical functions, which are typically limited in number. Figure LABEL:fig:lat_perc shows the loading latency distribution, revealing that 96.00% of context spaces can be loaded within one second. Loading latency correlates directly with context space size and I/O latency. When context spaces are cached, retrieval latency becomes negligible. Therefore, beyond our LRU replacement policy, larger context spaces (such as that for the settings app) can be permanently fixed in cache to further optimize performance.
@@
Table 5 presents CSAgent’s per task token consumption compared to the vanilla agent. The geometric mean of consumption overhead is 35.30%, with over 96.41% consisting of prompt tokens, which are less expensive than completion tokens. Generally, more complex tasks have proportionally lower overhead. For instance, AgentBench tasks have low average token consumption, so the large context space results in higher overhead. Conversely, AndroidWorld tasks are highly complex, requiring multimodal capabilities for screen recognition, leading to very high task token consumption and lower overhead. In comparison, runtime policy generation methods introduce 1.61× additional token consumption on AgentDojo (Shi et al., 2025a). CSAgent’s low token consumption results from development-phase policy generation that applies across all devices and users without regeneration, and from requiring only one additional LLM inference per task regardless of complexity (satisfying G2).